OpenStudy (anonymous):

What is an asymptote of the hyperbola given by the equation y^2/36 -x^2/1=1

OpenStudy (anonymous):

Answer choices: y=1/36x y=36x y=1/6x y=6x

OpenStudy (asnaseer):

have a look here: http://www.mathsisfun.com/geometry/hyperbola.html about three quarters of the way down that web page in section headed "Equation" it shows you how to get the asymptotes

OpenStudy (anonymous):

THANK YOU AGAIN!

OpenStudy (asnaseer):

yw :)